OpenStudy (anonymous):

6. what is the simplified form of the expression ? 5(14-2)2 —————— 2 ( 1point ) A . 60 b. 30 c. 72 d. 360 7. Evaluate u+xy ,if u = 18 , x = 10 , and y = 8. ( 1point ) A . 188 b.36 c. 98 d. 224

OpenStudy (igreen):

\(\dfrac{5(16-2)2}{2}\) Is the 2 all the way to the right on the top small and at the top? @fmkdfmr

OpenStudy (anonymous):

small

OpenStudy (igreen):

Okay. \(\dfrac{5(14-2)^2}{2}\) Subtract \(\dfrac{5(12)^2}{2}\) Exponent: \(\dfrac{12(144)}{2}\) Multiply: \(\dfrac{720}{2}\) \(360\)

OpenStudy (igreen):

\(u+xy\) \(u = 18 , x = 10 , y = 8\) \(18 + (10)(8)\) Multiply: \(18 + 80\) Add: \(98\)